OpenStudy (anonymous):

Find an equation of the tangent line to the graph of the function at the point (1, 1). y=x^(cosh(x)) y=????

OpenStudy (anonymous):

first you have to get the slope of the tangent line are you know how ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Do I get it by getting the derivative?

OpenStudy (anonymous):

yes right get the derivative then use the point (1,1)

OpenStudy (anonymous):

\[x^(\cosh(x)-1) * (\cosh(x)+xlog(x)\sinh(x))\] Do I just plug in the 1?

OpenStudy (anonymous):

are you sure about this result?

OpenStudy (anonymous):

I got it from WolframAlpha

OpenStudy (anonymous):

anyway a dont sure about this then plug in x=1 get y=?

OpenStudy (anonymous):

sorry get dy/dx =?

OpenStudy (anonymous):

I'm not sure, I'm really lost

OpenStudy (anonymous):

i got \[\frac{ dy }{ dx }=x ^{coshx}\left[ lnx *sinhx +\frac{ 1 }{ x }coshx\right]\]

OpenStudy (anonymous):

then \[\frac{ dy }{ dx }=1^{\cosh1}\left[ \ln1 *\sinh1 +\frac{ \cosh1 }{ 1 } \right]\]

OpenStudy (anonymous):

calculate it

OpenStudy (anonymous):

I get cosh(1)

OpenStudy (anonymous):

good then you have the data slope =cosh1 point=(1,1) can u use this data to create a line

OpenStudy (anonymous):

so the answer is cosh(x)? Because it's asking for a formula